OpenStudy (anonymous):

create an algebraic expression that can be simplified. example: x^2-5x+6/x^-9 simplify it

OpenStudy (luigi0210):

\[\frac{ x^2-5x+6 }{ x^-9 }=\frac{ (x-3)(x-2) }{ (x-3)(x+3) }\]

OpenStudy (luigi0210):

that's an x^2-9 at the bottom btw, my mistake

OpenStudy (anonymous):

no i need a knew one not the one i posted that was an example

OpenStudy (luigi0210):

Oh, does it need to have all different numbers?

OpenStudy (anonymous):

yea i think so

OpenStudy (anonymous):

it can be easy it just has to have varibles and needs to be able to be simplified

OpenStudy (luigi0210):

The previous? Or first?

OpenStudy (anonymous):

previous the answer

OpenStudy (anonymous):

ok im gonna turn it in ill tell u if your right

OpenStudy (anonymous):

she said we have to simplify and get a final answer @Luigi0210

OpenStudy (anonymous):

so its right so far were just not done

OpenStudy (luigi0210):

Well I showed it in the deleted post :l

OpenStudy (anonymous):

i know i copy and pasted she said we wernt done

OpenStudy (anonymous):

(x-2)(x-3)/(x-2)(x+2)=x^2-5x+6/x^2-4

OpenStudy (luigi0210):

\[\frac{ x-3 }{ x+2 }\]